# Quantifying the Dunkelflaute — An analysis of variable renewable energy droughts in Europe

### v3

@@ Abstract
Variable renewable energy droughts, also called “Dunkelflaute”, emerge as a challenge for climate-neutral energy systems based on variable renewables. Here we characterize European drought events for on- and offshore wind power, solar photovoltaics, and renewable technology portfolios, using 38 historic weather years and an advanced identification method. Their characteristics heavily depend on the chosen drought threshold, questioning the usefulness of single-threshold analyses. Applying a multi-threshold framework, we quantify how the complementarity of wind and solar power temporally and spatially alleviates drought frequency, return periods, duration, and severity within (portfolio effect) and across countries (balancing effect). We identify the most extreme droughts, which drive major discharging periods of long-duration storage in a fully renewable European energy system, based on a policy-relevant decarbonization scenario. Such events comprise sequences of shorter droughts of varying severity. The most extreme event occurred in winter 1996/97 and lasted 55 days in an idealized, perfectly interconnected setting. The average renewable availability during this period was still 47% of its long-run mean. System planners must consider such events when planning for storage and other flexibility technologies. Methodologically, we conclude that using arbitrary single calendar years is not suitable for modeling weather-resilient energy scenarios.
@@ 1 Introduction
Mitigating climate change requires deep reductions of greenhouse gas emissions that originate from fossil fuel use [lee_ipcc_2023]. Nuclear energy or fossil fuels in combination with carbon capture and storage are options for this [sepulveda_role_2018], but these firm low-carbon technologies face economic challenges and political controversies in many countries. Another major strategy for achieving climate neutrality is shifting energy supply to renewable energy sources, which are globally on the rise as they become increasingly cost-competitive [irena_renewable_2025]. However, the potential for dispatchable renewable energy sources such as hydro power, geothermal, or bioenergy is limited in most countries. Hence, future renewable energy systems are likely to heavily rely on variable wind and solar power [iea_net_2023, way_empirically_2022]. As these variable renewable energy (VRE) sources may become the most relevant primary energy sources in many countries [jacobson_100_2017, brown_response_2018, child_flexible_2019, pickering_diversity_2022, wang_accelerating_2023], energy systems are increasingly exposed to weather variability. In turn, spatial and temporal system flexibility is increasingly needed to match the supply of VRE sources with demand. This includes geographical balancing via transmission, different types of energy storage, and demand response [denholm_grid_2011, rasmussen_storage_2012, schlachtberger_benefits_2017, roth_geographical_2023, he_sector_2021, hunter_techno-economic_2021, tong_geophysical_2021].
@@
While future energy systems with high VRE shares have to deal with variability on different timescales, extreme periods of low wind and solar availability emerge as a key challenge for realizing such renewable energy systems. Often referred to as “Dunkelflaute”, these VRE drought periods are characterized by long-lasting and substantial shortages of wind and solar energy [raynaud_energy_2018, kittel_measuring_2024] and may cover large geographical areas. If the availability of firm low-carbon generation technologies is low, dealing with VRE drought events necessitates the use of long-duration storage as well as other temporal and spatial flexibility options [dowling_role_2020, sepulveda_design_2021]. Hence, understanding the spatial and temporal characteristics of VRE drought events is crucial for weather-resilient energy system planning and energy policy as well as for the design of energy markets and support instruments for generation and flexibility technologies. This includes questions as to how frequent, how long, and how severe such periods are, and assessing their spatial and temporal correlation across large-scale interconnected energy systems.
@@
Literature from different fields contributes to VRE drought analysis. Wind droughts in the United Kingdom are well-studied [cannon_using_2015, potisomporn_spatial_2022, potisomporn_evaluating_2023, abdelaziz_assessing_2024], focusing on frequency-duration distributions, return periods as well as spatial and temporal correlations of historic or future on- and offshore wind droughts. Similar analyses for wind power have been conducted for Ireland [leahy_persistence_2013], the North Sea [patlakas_low_2017], Germany [ohlendorf_frequency_2020], or, analyzing deviations from climatological means, globally [antonini_identification_2024]. Additionally, research interest in drought patterns of policy-relevant portfolios comprising wind and solar photovoltaics (PV) is growing. Historic, future, and synthetic weather data of various world regions have been analyzed, such as Europe [raynaud_energy_2018, kies_critical_2021, kapica_potential_2024, hu_implications_2023, breyer_reflecting_2022, van_der_most_temporally_2024, van_duinen_meteorological_2025], the U.S. [rinaldi_wind_2021, bracken_standardized_2024, bracken_seasonal_2025], China [li_renewable_2024, lei_frequency_2024], India [gangopadhyay_role_2022], Germany [kaspar_climatological_2019, mockert_meteorological_2023], Hungary [mayer_probabilistic_2023], Japan [ohba_climatology_2022], and Australia [richardson_climate_2023]. Besides regional and seasonal variations for single technologies and VRE portfolios, a general finding is that combining wind and solar within regions, as well as considering balancing across regions, can mitigate drought characteristics.
@@
Different methodological approaches have been used for renewable drought identification [ohlendorf_frequency_2020, potisomporn_extreme_2024, kittel_measuring_2024, li_renewable_2024, biewald_evaluation_2025, wilczak_wind_2025]. Droughts can be defined as periods of consecutive time steps with renewable availability below a certain drought threshold, either with fixed [raynaud_energy_2018, rinaldi_wind_2021, kies_critical_2021, gangopadhyay_role_2022, kapica_potential_2024, kaspar_climatological_2019, ohba_climatology_2022] or variable duration [potisomporn_extreme_2024], using various identification methods. For instance, these periods can be identified by searching for an availability constantly below the threshold [potisomporn_spatial_2022, potisomporn_evaluating_2023, abdelaziz_assessing_2024, ohlendorf_frequency_2020, kies_critical_2021, breyer_reflecting_2022, kaspar_climatological_2019, mayer_probabilistic_2023, ohba_climatology_2022], a mean availability over a certain averaging interval below the threshold [abdelaziz_assessing_2024, leahy_persistence_2013, patlakas_low_2017, ohlendorf_frequency_2020, raynaud_energy_2018, rinaldi_wind_2021, gangopadhyay_role_2022, kapica_potential_2024, mockert_meteorological_2023], or the deviation of the availability from a drought threshold [potisomporn_extreme_2024] or its climatological mean [antonini_identification_2024, stoop_climatological_2024]. Drought thresholds are either exogenously and presumably arbitrarily set [cannon_using_2015, potisomporn_extreme_2024, abdelaziz_assessing_2024, leahy_persistence_2013, patlakas_low_2017, ohlendorf_frequency_2020, kies_critical_2021, kaspar_climatological_2019, mockert_meteorological_2023, mayer_probabilistic_2023, ohba_climatology_2022], or derived from the data analyzed, such as a fraction of the time-invariant mean [raynaud_energy_2018, rinaldi_wind_2021, kapica_potential_2024] or maximum [breyer_reflecting_2022] availability, or of its time-variant climatological mean [gangopadhyay_role_2022, antonini_identification_2024, stoop_climatological_2024]. Other researchers have proposed using copulas to construct joint cumulative distribution functions of drought severity and duration [otero_copula-based_2022] or standardized indices to monitor and compare drought events [allen_standardised_2023]. As each of these approaches has strengths and weaknesses, no standard drought identification method has yet emerged in the literature [kittel_measuring_2024].
@@
Here we analyze and compare VRE drought events for single renewable technologies and for a fully renewable European electricity system, using VRE portfolio assumptions from a policy-relevant decarbonization scenario from European transmission grid operators. We do so for 33 individual European countries (EU27, the United Kingdom, Norway, Switzerland, and the Western Balkans), and for a pan-European “copperplate” scenario with perfect interconnection across all countries that allows for unconstrained geographical balancing. The latter represents an idealized scenario that allows identifying the upper bound of the value of geographical balancing for mitigating renewable energy droughts. We draw on a large dataset of historical VRE availability factor time series covering 38 years [de_felice_entso-e_2022]. We use an advanced, open-source algorithm that fully captures unique drought periods, properly accounts for brief periods of higher renewable availability within longer drought events, and avoids arbitrary threshold choices [kittel_measuring_2024]. Note that the question what mix of storage and other flexibility technologies optimally facilitates weather-resilient renewable energy system is not in the focus of our analysis. This has to be investigated using computationally intensive numerical energy system models. However, the methods and findings presented here can help to identify relevant weather years that should be used in such model analyses.
@@ 2 Results
We define a renewable energy drought as a period during which the moving average of the hourly VRE availability factor, also referred to as capacity factor, is below a given qualification threshold [kittel_measuring_2024]. The search algorithm used here iteratively decreases the moving average window, which initially identifies the longest-lasting droughts, followed by progressively shorter ones. As there is no consensus in the literature on appropriate drought thresholds, we apply a large number of thresholds to each investigated region and technology, ranging in 5% increments from 10% to 100% of the mean renewable availability across all investigated years [kittel_measuring_2024]. While low thresholds identify severe droughts, higher thresholds increasingly include more moderate drought events. By scaling drought thresholds relative to the mean availability factor of all hours in the data of each region-technology setting (henceforth referred to as “relative thresholds”), we account for regional and technology-specific differences in generation potentials. This enables meaningful cross-regional and cross-technological comparisons. Section 4.1 details the renewable energy drought definition and identification method. To give an example, an identified wind power drought with a 24-hour duration at a threshold $\tau_{0.1}$ should be interpreted such that less than 10% of the long-run average electricity generation of wind power is available on average during a 24-hour period.
In the main part of the paper, we illustrate selected results for a limited set of countries. The SI provides additional results and information on the thresholds used. Interactive, high-resolution versions of all graphs that allow for retrieving concrete values for each of the illustrated 3D-plots as well as an additional animated graph are available online [kittel_high_2025].
@@ 2.1 Complex renewable energy drought dynamics and complementarities across regions and technologies
Figure 1: Drought patterns in 1996 and 1997 for selected regions.
@@
In contrast, severe long-lasting onshore and offshore wind droughts can occur throughout the year but tend to be more frequent in summer (Figure 1, panels c and d), which is in line with general wind seasonality [kaspar_climatological_2019]. The multi-threshold perspective further reveals that severe wind droughts identified by lower thresholds can be sequential and may occur within contiguous below-average wind periods detected at higher thresholds. The latter may last up to several months. While such long-lasting events usually occur in summer (compare Figures SI.3-SI.5), they may also take place in winter in some countries (e.g., in 1996/97, Figure 1). On- and offshore droughts are generally correlated, but may differ in severity (Figure 1, panels c and d).
@@ 2.2 Frequency-duration distribution: short drought events occur much more often than longer ones
Cumulative frequency-duration distributions of drought events across all investigated relative drought thresholds $\tau_{i}$ with $i\in[0.1,...,1]$, sorting the yearly frequencies of all events that are at least as long as a given duration. White space indicates the absence of droughts for given thresholds in the data. The contour lines represent the threshold-specific yearly frequency. For illustration, the distributions are truncated at 360 hours, i.e., they show events with a maximum duration of just above two weeks. Yearly frequencies of events lasting at least two days, one week, and a fortnight are marked for a relative threshold of $\tau_{0.75}$. An interactive version of this Figure which allows zooming and rotating is available online [kittel_high_2025].
@@
The cumulative frequency-duration distribution substantially differs between wind and solar power. For PV and low thresholds, we identify many events that last less than one day. They reflect the typical diurnal fluctuations of solar PV, with zero production at night-time [ueckerdt_analyzing_2015]. Increasing thresholds lead to fewer PV drought events, as many night-time periods of zero electricity generation merge into multi-day events. When the threshold approaches $\tau_{1.0}$, the identified droughts increasingly reflect a strong solar seasonality in Europe, with lower availability in winter than in summer. For on- and offshore wind power, droughts lasting less than one day occur less often. This is because wind power does not have regular non-availability at night-time as solar PV, but fluctuates less regularly. In turn, events that last longer than one day are more frequent, reflecting a typical multi-day variability of wind power [apt_spectrum_2007, ueckerdt_analyzing_2015].
@@
Figure 2 also illustrates the geographical balancing effect, as droughts occur less often in an ideal European interconnection than in single countries. This is true for both single VRE technologies and portfolios. The reason for this is that periods of low solar and wind availability are not perfectly correlated across Europe [roth_geographical_2023]. For example, assuming a relative threshold $\tau_{0.75}$, there have been on average approximately eight and five VRE portfolio droughts per year in Germany and Spain that lasted at least one week, respectively, while less than four of such events occurred in Europe. For on- and offshore wind, the balancing effect mitigates drought frequency, particularly for drought events identified by lower thresholds, i.e., more severe wind droughts. This effect is visualized by the additional white space in Figure 2 and the frequency-duration distributions in Figure SI.7 when in the European copperplate scenario compared to Germany or Spain, respectively.
@@ 2.5 Portfolio and balancing effects mitigate maximum drought durations
Comparing weighted averages over all thresholds and countries, the maximum drought duration of a VRE portfolio decreases by 64%, 52%, or 47% compared to solar PV, onshore wind, or offshore wind alone (Figure 6). This portfolio effect is more pronounced for PV than for wind power, as onshore and offshore wind power complement both diurnal and seasonal shortages of PV. In contrast, wind power does not incur diurnal shortages. Assuming a European copperplate (CP), the portfolio effect (-80%) for solar PV is even stronger than for any isolated country (Table SI.1). This is because in a larger geographical area, there is more complementary wind power available to compensate for solar PV shortages [roth_geographical_2023]. In turn, the portfolio effect is less pronounced for onshore wind power in the European interconnection (-45%) as PV availability is more homogeneously distributed. For offshore wind power, the portfolio effect is higher (-70%), as only a subset of countries features this technology, such that solar and wind onshore resources of all countries can compensate for extreme offshore droughts in a subset of countries.
@@
Previous studies that use threshold-based identification methods typically use only a limited number of thresholds (e.g., [leahy_persistence_2013, cannon_using_2015, patlakas_low_2017, kaspar_climatological_2019, ohlendorf_frequency_2020, kies_critical_2021, potisomporn_extreme_2024, potisomporn_extreme_2024, abdelaziz_assessing_2024, mockert_meteorological_2023, mayer_probabilistic_2023, ohba_climatology_2022]). However, the findings presented above indicate that drought characteristics strongly depend on the chosen threshold. To account for this, we introduce a multi-threshold “drought mass” metric to identify the most extreme portfolio drought events and show that it identifies events that are relevant for long-duration electricity storage in renewable European energy systems. In contrast to previous single-threshold approaches, the drought mass quantifies extreme events by integrating both drought duration and severity across a wide range of thresholds, thus properly accounting for the complex and diverse renewable drought patterns discussed in Section 2.1. It accumulates the number of drought hours identified at thresholds $\tau_{i\leq 0.75}$ that occur within contiguous events identified at $\tau_{0.75}$, i.e., excluding droughts found by very high thresholds ($\tau_{i>0.75}$). The event with the highest score within a each pair of consecutive years in the data (i.e., 1982-1983, 1983-1984, 1984-1985, etc.) is identified as the most extreme drought event within this time frame. Section 4.2 further elaborates on the drought mass mechanics.

### v4

# Multi-threshold time series analysis enables characterization of variable renewable energy droughts in Europe

@@ Abstract
Variable renewable energy droughts, so called Dunkelflaute events, emerge as a challenge for climate-neutral energy systems based on variable renewables. Here we characterize European drought events for on- and offshore wind power, solar photovoltaics, and renewable technology portfolios, using 38 historic weather years and an advanced identification method. Their characteristics heavily depend on the chosen drought threshold, questioning the usefulness of single-threshold analyses. Applying a multi-threshold framework, we quantify how the complementarity of wind and solar power temporally and spatially alleviates drought frequency, return periods, duration, and severity within (portfolio effect) and across countries (balancing effect). We identify the most extreme droughts, which drive major discharging periods of long-duration storage in a fully renewable European energy system, based on a policy-relevant decarbonization scenario. Such events comprise sequences of shorter droughts of varying severity. The most extreme event occurred in winter 1996/97 and lasted 55 days in an idealized, perfectly interconnected setting. The average renewable availability during this period was still 47% of its long-run mean. System planners must consider such events when planning for storage and other flexibility technologies. Methodologically, we conclude that using arbitrary single calendar years is not suitable for modeling weather-resilient energy scenarios.
@@ 1 Introduction
Mitigating climate change requires deep reductions of greenhouse gas emissions that originate from fossil fuel use [13]. Nuclear energy or fossil fuels in combination with carbon capture and storage are options for this [88], but these firm low-carbon technologies face economic challenges and political controversies in many countries. Another major strategy for achieving climate neutrality is shifting energy supply to renewable energy sources, which are globally on the rise as they become increasingly cost-competitive [42]. However, the potential for dispatchable renewable energy sources such as hydro power, geothermal, or bioenergy is limited in most countries. Hence, future renewable energy systems are likely to heavily rely on variable wind and solar power [41, 99]. As these variable renewable energy (VRE) sources may become the most relevant primary energy sources in many countries [43, 10, 15, 67, 98], energy systems are increasingly exposed to weather variability. In turn, spatial and temporal system flexibility is increasingly needed to match the supply of VRE sources with demand. This includes geographical balancing via transmission, different types of energy storage, and demand response [19, 72, 85, 77, 36, 40, 94].
@@
While future energy systems with high VRE shares have to deal with variability on different timescales, extreme periods of low wind and solar availability emerge as a key challenge for realizing such renewable energy systems. Often referred to as “Dunkelflaute”, these VRE drought periods are characterized by long-lasting and substantial shortages of wind and solar energy [73, 51] and may cover large geographical areas. If the availability of firm low-carbon generation technologies is low, dealing with VRE drought events necessitates the use of long-duration storage as well as other temporal and spatial flexibility options [21, 89]. Hence, understanding the spatial and temporal characteristics of VRE drought events is crucial for weather-resilient energy system planning and energy policy as well as for the design of energy markets and support instruments for generation and flexibility technologies. This includes questions as to how frequent, how long, and how severe such periods are, and assessing their spatial and temporal correlation across large-scale interconnected energy systems.
@@
Literature from different fields contributes to VRE drought analysis. Wind droughts in the United Kingdom are well-studied [14, 70, 68, 1], focusing on frequency-duration distributions, return periods as well as spatial and temporal correlations of historic or future on- and offshore wind droughts. Similar analyses for wind power have been conducted for Ireland [53], the North Sea [64], Germany [62], or, analyzing deviations from climatological means, globally [3]. Additionally, research interest in drought patterns of policy-relevant portfolios comprising wind and solar photovoltaics (PV) is growing. Historic, future, and synthetic weather data of various world regions have been analyzed, such as Europe [73, 47, 45, 39, 9, 96, 97], the U.S. [75, 7, 8], China [55, 54], India [29], Germany [46, 58], Hungary [57], Japan [61], and Australia [74]. Besides regional and seasonal variations for single technologies and VRE portfolios, a general finding is that combining wind and solar within regions, as well as considering balancing across regions, can mitigate drought characteristics.
@@
Different methodological approaches have been used for renewable drought identification [62, 69, 51, 55, 6, 100]. Droughts can be defined as periods of consecutive time steps with renewable availability below a certain drought threshold, either with fixed [73, 75, 47, 29, 45, 46, 61] or variable duration [69], using various identification methods. For instance, these periods can be identified by searching for an availability constantly below the threshold [70, 68, 1, 62, 47, 9, 46, 57, 61], a mean availability over a certain averaging interval below the threshold [1, 53, 64, 62, 73, 75, 29, 45, 58], or the deviation of the availability from a drought threshold [69] or its climatological mean [3, 91]. Drought thresholds are either exogenously and presumably arbitrarily set [14, 69, 1, 53, 64, 62, 47, 46, 58, 57, 61], or derived from the data analyzed, such as a fraction of the time-invariant mean [73, 75, 45] or maximum [9] availability, or of its time-variant climatological mean [29, 3, 91]. Other researchers have proposed using copulas to construct joint cumulative distribution functions of drought severity and duration [63] or standardized indices to monitor and compare drought events [2]. As each of these approaches has strengths and weaknesses, no standard drought identification method has yet emerged in the literature [51].
@@
Here we analyze and compare VRE drought events for single renewable technologies and for a fully renewable European electricity system, using VRE portfolio assumptions from a policy-relevant decarbonization scenario from European transmission grid operators. We do so for 33 individual European countries (EU27, the United Kingdom, Norway, Switzerland, and the Western Balkans), and for a pan-European “copperplate” scenario with perfect interconnection across all countries that allows for unconstrained geographical balancing. The latter represents an idealized scenario that allows identifying the upper bound of the value of geographical balancing for mitigating renewable energy droughts. We draw on a large dataset of historical VRE availability factor time series covering 38 years [18]. We use an advanced, open-source algorithm that fully captures unique drought periods, properly accounts for brief periods of higher renewable availability within longer drought events, and avoids arbitrary threshold choices [51]. Note that the question what mix of storage and other flexibility technologies optimally facilitates weather-resilient renewable energy system is not in the focus of our analysis. This has to be investigated using computationally intensive numerical energy system models. However, the methods and findings presented here can help to identify relevant weather years that should be used in such model analyses.
@@ 2 Results
We define a renewable energy drought as a period during which the moving average of the hourly VRE availability factor, also referred to as capacity factor, is below a given qualification threshold [51]. The search algorithm used here iteratively decreases the moving average window, which initially identifies the longest-lasting droughts, followed by progressively shorter ones. As there is no consensus in the literature on appropriate drought thresholds, we apply a large number of thresholds to each investigated region and technology, ranging in 5% increments from 10% to 100% of the mean renewable availability across all investigated years [51]. While low thresholds identify severe droughts, higher thresholds increasingly include more moderate drought events. By scaling drought thresholds relative to the mean availability factor of all hours in the data of each region-technology setting (henceforth referred to as “relative thresholds”), we account for regional and technology-specific differences in generation potentials. This enables meaningful cross-regional and cross-technological comparisons. Section 4.1 details the renewable energy drought definition and identification method. To give an example, an identified wind power drought with a 24-hour duration at a threshold $\tau_{0.1}$ should be interpreted such that less than 10% of the long-run average electricity generation of wind power is available on average during a 24-hour period.
In the main part of the paper, we illustrate selected results for a limited set of countries. The SI provides additional results and information on the thresholds used. Interactive, high-resolution versions of all graphs that allow for retrieving concrete values for each of the illustrated 3D-plots as well as an additional animated graph are available online [52].
@@ 2.1 Complex renewable energy drought dynamics and complementarities across regions and technologies
Figure 1: Drought patterns in 1996 and 1997 for selected regions, corresponding to the following technology settings: a) renewable technology portfolio combining solar PV as well as on- and offshore wind, b) solar PV, c) onshore wind, and d) offshore wind.
@@
In contrast, severe long-lasting onshore and offshore wind droughts can occur throughout the year but tend to be more frequent in summer (Figure 1, panels c and d), which is in line with general wind seasonality [46]. The multi-threshold perspective further reveals that severe wind droughts identified by lower thresholds can be sequential and may occur within contiguous below-average wind periods detected at higher thresholds. The latter may last up to several months. While such long-lasting events usually occur in summer (compare Figures SI.3-SI.5), they may also take place in winter in some countries (e.g., in 1996/97, Figure 1). On- and offshore droughts are generally correlated, but may differ in severity (Figure 1, panels c and d).
@@ 2.2 Frequency-duration distribution: short drought events occur much more often than longer ones
Cumulative frequency-duration distributions of drought events across all investigated relative drought thresholds $\tau_{i}$ with $i\in[0.1,...,1]$, sorting the yearly frequencies of all events that are at least as long as a given duration. White space indicates the absence of droughts for given thresholds in the data. The contour lines represent the threshold-specific yearly frequency. For illustration, the distributions are truncated at 360 hours, i.e., they show events with a maximum duration of just above two weeks. Yearly frequencies of events lasting at least two days, one week, and a fortnight are marked for a relative threshold of $\tau_{0.75}$. An interactive version of this Figure which allows zooming and rotating is available online [52].
@@
The cumulative frequency-duration distribution substantially differs between wind and solar power. For PV and low thresholds, we identify many events that last less than one day. They reflect the typical diurnal fluctuations of solar PV, with zero production at night-time [95]. Increasing thresholds lead to fewer PV drought events, as many night-time periods of zero electricity generation merge into multi-day events. When the threshold approaches $\tau_{1.0}$, the identified droughts increasingly reflect a strong solar seasonality in Europe, with lower availability in winter than in summer. For on- and offshore wind power, droughts lasting less than one day occur less often. This is because wind power does not have regular non-availability at night-time as solar PV, but fluctuates less regularly. In turn, events that last longer than one day are more frequent, reflecting a typical multi-day variability of wind power [4, 95].
@@
Figure 2 also illustrates the geographical balancing effect, as droughts occur less often in an ideal European interconnection than in single countries. This is true for both single VRE technologies and portfolios. The reason for this is that periods of low solar and wind availability are not perfectly correlated across Europe [77]. For example, assuming a relative threshold $\tau_{0.75}$, there have been on average approximately eight and five VRE portfolio droughts per year in Germany and Spain that lasted at least one week, respectively, while less than four of such events occurred in Europe. For on- and offshore wind, the balancing effect mitigates drought frequency, particularly for drought events identified by lower thresholds, i.e., more severe wind droughts. This effect is visualized by the additional white space in Figure 2 and the frequency-duration distributions in Figure SI.7 when in the European copperplate scenario compared to Germany or Spain, respectively.
@@ 2.5 Portfolio and balancing effects mitigate maximum drought durations
Comparing weighted averages over all thresholds and countries, the maximum drought duration of a VRE portfolio decreases by 64%, 52%, or 47% compared to solar PV, onshore wind, or offshore wind alone (Figure 6). This portfolio effect is more pronounced for PV than for wind power, as onshore and offshore wind power complement both diurnal and seasonal shortages of PV. In contrast, wind power does not incur diurnal shortages. Assuming a European copperplate (CP), the portfolio effect (-80%) for solar PV is even stronger than for any isolated country (Table SI.1). This is because in a larger geographical area, there is more complementary wind power available to compensate for solar PV shortages [77]. In turn, the portfolio effect is less pronounced for onshore wind power in the European interconnection (-45%) as PV availability is more homogeneously distributed. For offshore wind power, the portfolio effect is higher (-70%), as only a subset of countries features this technology, such that solar and wind onshore resources of all countries can compensate for extreme offshore droughts in a subset of countries.
@@
Previous studies that use threshold-based identification methods typically use only a limited number of thresholds (e.g., [53, 14, 64, 46, 62, 47, 69, 69, 1, 58, 57, 61]). However, the findings presented above indicate that drought characteristics strongly depend on the chosen threshold. To account for this, we introduce a multi-threshold “drought mass” metric to identify the most extreme portfolio drought events and show that it identifies events that are relevant for long-duration electricity storage in renewable European energy systems. In contrast to previous single-threshold approaches, the drought mass quantifies extreme events by integrating both drought duration and severity across a wide range of thresholds, thus properly accounting for the complex and diverse renewable drought patterns discussed in Section 2.1. It accumulates the number of drought hours identified at thresholds $\tau_{i\leq 0.75}$ that occur within contiguous events identified at $\tau_{0.75}$, i.e., excluding droughts found by very high thresholds ($\tau_{i>0.75}$). The event with the highest score within a each pair of consecutive years in the data (i.e., 1982-1983, 1983-1984, 1984-1985, etc.) is identified as the most extreme drought event within this time frame. Section 4.2 further elaborates on the drought mass mechanics.